# Bilevel subsidy-enabled mobility hub network design with perturbed utility coalitional choice-based assignment

### v1

@@ 3 Proposed mobility hub platform design methodology
We first present the multimodal network structure and service assumptions. We then present the model formulations in detail. The model formulation for MH platform design takes the general stochastic MaaS assignment game from (Liu et al.,, 2024) and re-formulates it within an equivalent PURC framework. The goal is to provide a scalable framework that jointly decides the service decisions from the operator side and path choices from the customer side, while the platform sets service subsidies.
3.1 Network structure and model assumption
@@
The general notation for MaaS platform design used for the MH platform design follows a multicommodity flow network design problem. The multimodal network $G$, denoted by $(N,A)$, consists of all nodes $N$ and links $A$, serving all origin-destination (OD) groups $S$. Origins and destinations are aggregated into centroids $N_{S}\subset N$. The services include a single fixed route transit (FT) operator acting as the platform for designing the MHs, and a set of MOD services that can choose to participate as feeders for these MHs. The model can be trivially modified to include fixed route buses run by other operators as feeders as well, but for simplicity in notation we assume this MH system deals only with a single fixed route platform/operator and one or more MOD potential feeder operators. The FT operator $f$ belongs to a singleton set $F$ and provides service on its own subnetwork $G_{f}$, which consists of node set $N_{f}\subseteq N_{F}\subseteq N$ and link set $A_{f}\subseteq A_{F}\subseteq A$. Similarly, each MOD fleet $m$ belongs to the MOD operator set $M$, with each operating a subnetwork $G_{m}$ consisting of node set $N_{m}\subseteq N_{M}\subseteq N$ and service link set $A_{m}\subseteq A_{M}\subseteq A$. A MOD service link represents a pickup and origin and drop-off at destination, where the link attributes reflect average performance of the fleet serving this request. Congestion effects, such as waiting for service, are captured by a node capacity at the origin. These ”in-platform” links are store-and-forward links with service queues. Meanwhile, each MOD operator $m$ decides on the node capacities $z_{i}$ for all nodes $i$ belonging to node set $N_{m}$.
@@
A dummy subnetwork $G_{O}$ is used to represent travel outside of the MH platform’s multimodal services (e.g., private car, telecommuting, or an alternative platform), i.e. ”out-of-platform” links. A set of transfer links $A_{0}$ are used to connect centroids $N_{S}$ with service nodes $N_{F}\cup N_{M}\cup N_{O}$, as well as connections between layers of subnetworks to construct the overall multimodal network. Specifically, links connecting other layers to a MOD subnetwork layer are called MOD access links. Fig. 1 illustrates the multimodal network structure. Travelers choose their preferred paths encompassing combinations of services across the network for each OD pair $s\in S$, which may use the platform to access the hub and taken the fixed route transit to the destination (and vice versa), use their own modes to get to the transit station to go to the destination, or use their own mode entirely to get to the destination.
@@
In the case of a mobility hub-oriented system, we consider a platform operator, which may or may not be a leading fixed-route transit operator, working with MOD operators to provide services at designated MHs through a subset of nodes $H\subset N$. These operators may be micromobility providers, ridehail, or microtransit services. The hubs in this setting represent cyberphysical gateways in which participating MOD operators that pick up or drop off passengers purchasing trip bundles within these geofenced hubs could potentially be subsidized. Potential MOD feeder services are introduced that are centered around each hub $h\in H$. Direct access between hubs and surrounding service nodes is represented by a set of added links $A_{H}$. The total flow of all direct links $A_{h}$ leading to hub $h$ shall not exceed the capacity of the hub.
Figure 1: Network illustration
@@
As shown in Fig. 1, we separate subnetworks to be ”in-platform” and ”out-of-platform”. We focus on the service strategy decisions of operators in-platform, including link-based prices and service capacities. Since out-of-platform includes all other options out there (including driving, walking, telecommuting), they are assumed to be uncapacitated and we do not explicitly model competition between the designed platform and other external modes. Instead, we focus on the cooperative game between operators in-platform.
@@
Inside the platform, the FT operator acts as the regulator of a MH platform that engages with MOD services to connect travelers to the FT subnetwork. Therefore, only the price of the MOD service links in the in-platform subnetworks are part of the decision variables. Other decision variables include the service capacities at either the MH or the MOD service nodes within the in-platform subnetworks. Table 1 provides the full list of notations. In addition, we list the model assumptions as follows. To avoid confusion, we define travel cost as the non-monetary penalties associated with links such as travel time. Service price is the monetary values required for using each link. Both terms are converted to utilities.
@@
$A_{0}$ Transfer links between centroids and service nodes, and between layers
@@
$A_{H}$ Added links connecting hubs and service nodes
$A_{h}$ Direct links leading to hub $h\in H$
@@
$z_{i}$ Capacity of MOD node $i\in N_{M}$
@@
$a_{i,l}$ Node-link incidence: 1-1 if origin, +1+1 if destination, 0 otherwise
$\hat{p}_{l}$ Cap of service price for link $l\in A$
@@
$\alpha$ Weight of operator utility
@@
$x_{s,l}$ Normalized flow on link $l\in A$ for OD pair $s\in S$
$v_{i}$ Proportion of maximum capacity opened for MOD node $i\in N_{M}$
$p_{l}$ Service price on link $l\in A$; if lower than operator’s price, reflects a subsidy from the platform
@@
    Travel cost of the FT service links and out-of-platform links are fixed.
@@
    Transfer links between centroids and FT nodes have fixed travel cost including wait and walk time, which represents the overall cost of accessing FT services.
@@
    MOD access links are store-and-forward links with service queues, meaning that the access flow remains uncongested until capacity is reached. Once capacity is exceeded, excess flows are diverted to other links with available capacity.
@@
    All FT service links and links in the out-of-service subnetwork have fixed travel cost and are uncapacitated. Congestion from background traffic are directly incorporated into the travel costs. If an out-of-service link reflects multiple modes, its travel cost reflects the expected minimized disutility option of those modes. Alternatively each parallel mode may be modeled separately.
@@
    Operating cost of MOD services consists of service link costs and access capacity allocation costs.
3.2 Lower-level traveler-operator coalition choice model using PURC based structure
The MH platform design problem is adapted from the stochastic MaaS assignment game model in Liu et al., (2024). The assignment game is decomposed into a bilevel problem where the upper level involves platform decisions, such as pricing and subsidies, while the lower level involves the joint decisions of the travelers and mobility operators. In the adaptation to the MH platform design problem, the platform is also the FT operator making decisions about how much to subsidize the MOD operators in the upper level, while the lower level determines the MH and MOD capacities and traveler flows.
Instead of using the path-based SUE-style formulation in Liu et al., (2024), we adapt the link-based PURC model proposed by Fosgerau et al., (2022) to capture the probabilistic matching mechanism. Eq. (1) is the link-based stochastic assignment problem $L_{1}$ based on the PURC concept for a MH platform.
$L_{1}\min_{\{x_{l,s},\,v_{i},\}}\quad$ $\Phi_{1}=\sum_{l\in A}\sum_{s\in S}\bar{q}d_{l}x_{l,s}^{2}+\sum_{l\in A}\sum_{s\in S}\bar{q}d_{l}\left(p_{l}+c^{t}_{l}\right)x_{l,s}+\alpha\left(\sum_{l\in A}\sum_{s\in S}d_{l}c^{o}_{l}x_{l,s}q_{s}+\sum_{i\in N_{M}}z_{i}c_{i}v_{i}\right)$ (1a)
s.t. $\sum_{l\in A}a_{i,l}x_{l,s}=\begin{cases}-1,\quad\forall i=o,\\ 1,\quad\forall i=d,\\ 0,\quad\forall i\in N\setminus\{o,d\}\end{cases}\quad\forall s=(o,d)\in S,\,\forall i\in N$ (1b)
$\sum_{l\in A^{+}_{0,i}}\sum_{s\in S}x_{l,s}q_{s}\leq z_{i}v_{i}\quad\forall i\in N_{M}$ (1c)
$\sum_{l\in A^{+}_{i,h}}\sum_{s\in S}x_{l,s}q_{s}\leq z_{i}v_{i}\quad\forall h\in N_{H}$ (1d)
@@ 3.2 Lower-level traveler-operator coalition choice model using PURC based structure
The decision variables are the link flow assignment $x_{l,s}$ and MOD node capacity $v_{i}$. The service price $p_{l}$ is the decision variable in the upper level model, which is presented in the following subsection. In $\Phi_{1}$, $p_{l}$ is treated as an input. Eq. (1b) ensures the link flow conservation. For each OD pair $s=(o,d)$ with $o$ being the origin and $d$ being the destination, the source node $o$ initiates all normalized flow, and the sink node $d$ terminates all normalized flow. Eq. (1c) controls the capacity and congestion effect of the platform service node. All access links flowing into the service nodes are bounded by the assigned capacity the operator determines. Eq. (1d) is dedicated to mobility hub facilities. For all service flows ending at mobility hubs, the sum of flow needs to be capped by the assigned mobility hub capacity. Eqs. (1e - 1f) are the bounds for the two decision variables.
The objective consists of three terms: the perturbed utility, traveler utility, and the operator utility. Introduced by Fosgerau et al., (2022), the PURC model follows the general RUM structure with link-based property to capture the dispersion of flows across a network. The flow assignment variable $x_{l,s}$ represents the probability of choosing link $l$ for travelers in OD pair $s$. The first term in Eq. (1a) depicts the dispersion effect. The second term is the sum of traveler utility, which depicts the sum of all link utilities. We simply the term by only involving a general cost element and a service price element. It can be expanded to involve all utility related elements such as externalities, as long as they can be converted to utility values. For a comprehensive explanation of the disperse effect and the utility features that can be included, readers can refer to Fosgerau et al., (2022) for more information. The original PURC objective is constructed by combining the first and second terms. We show this in Proposition 1.
@@
The first and second terms combined in Eq. (1a) are equivalent to the PURC objective proposed by Fosgerau et al., (2022).
@@
$\min_{\{x_{l,s},\,v_{i},\}}\quad\Phi_{1}=\bar{q}\sum_{s\in S}d_{l}(x_{l,s}^{2}+\left(p_{l}+c^{t}_{l}\right)x_{l,s})$ (2)
The coefficient $\bar{q}$ is a constant. The quadratic term is the perturbed term $F(x)$, and the linear term is the link-additive utility term $U(x)$. Both terms are weighted by traversed link length $d_{l}$. The structure is therefore identical to the PURC objective proposed in Fosgerau et al., (2022) by multiplying with a constant $\bar{q}$. ∎
We introduce the third term to form the coalition between travelers and operators in a similar fashion to the model proposed in chapter 4 of Liu, (2024). In Proposition 2, we show that $\Phi_{1}$ provides a flow assignment model with a traveler-operator coalition under the rule of PURC.
@@
We divide Eq. (1a) by the constant term $\sum{q_{s}}/{|S|}$ and write it as Eq. (3).
$\min_{\{x_{l,s},\,v_{i},\}}\quad\Phi^{\prime}_{1}=\sum_{s\in S}\sum_{l\in A}d_{l}x_{l,s}^{2}+\sum_{s\in S}\sum_{l\in A}d_{l}\left(p_{l}+c^{t}_{l}+\alpha c^{o}_{l}\frac{q_{s}}{\bar{q}}\right)x_{l,s}+\alpha\left(\sum_{i\in N_{M}}\frac{z_{i}c_{i}v_{i}}{\bar{q}}\right)$ (3)
@@
$\min_{\{x_{l,s},\,v_{i},\}}\quad L^{\prime}_{1}$ $=\sum_{s\in S}\sum_{l\in A}d_{l}x_{l,s}^{2}+\sum_{s\in S}\sum_{l\in A}d_{l}\left(p_{l}+c^{t}_{l}+\alpha c^{o}_{l}\frac{q_{s}}{\bar{q}}\right)x_{l,s}+\alpha\left(\sum_{i\in N_{M}}\frac{z_{i}c_{i}v_{i}}{\bar{q}}\right)$ (4a)
$+\sum_{s\in S}\sum_{i\in N}\mu_{i,s}(\sum_{l}a_{i,l}x_{l,s}-f_{i,s})$ (4b)
$+\sum_{i\in N_{M}}\lambda_{i}(\sum_{l\in A^{-}_{0,i}}\sum_{s\in S}x_{l,s}q_{s}-z_{i}v_{i})$ (4c)
$+\sum_{l\in A}\sum_{s\in S}\beta_{l,s}(x_{l,s}-1)+\sum_{i\in N_{M}}\pi_{i}(v_{i}-1)$ (4d)
@@
$\frac{\partial L^{\prime}_{1}}{x_{l,s}}$ $=2d_{l}x_{l,s}+d_{l}\left(p_{l}+c^{t}_{l}+\alpha\frac{c^{o}_{l}q_{s}}{\bar{q}}\right)$ (5a)
$+\sum_{i\in{l^{+},l^{-}}}\mu_{i,s}a_{i,l}+\sum_{i\in{l^{+},l^{-}}}\lambda_{i}q_{s}+\beta_{l,s}=0$ (5b)
Because the chosen links will have strictly positive flows, the $v_{i}$ connecting those links also fulfill the optimality conditions.
$\frac{\partial L^{\prime}_{1}}{v_{i}}$ $=\alpha\frac{z_{i}c_{i}}{\bar{q}}-\lambda_{i}z_{i}=0$ (6a)
By re-writing Eq. (6), we obtain $\lambda_{i}=\frac{\alpha c_{i}}{\bar{q}}$. By plugging it in Eq. (5), we have Eq. (7).
$\frac{\partial L^{\prime}_{1}}{x_{l,s}}$ $=2d_{l}x_{l,s}+d_{l}\left(p_{l}+c^{t}_{l}+\alpha\frac{c^{o}_{l}q_{s}}{\bar{q}}\right)$ (7a)
$+\sum_{i\in{l^{+},l^{-}}}\mu_{i,s}a_{i,l}+\sum_{i\in{l^{+},l^{-}}}\alpha c_{i}\frac{q_{s}}{\bar{q}}+\beta_{l,s}=0$ (7b)
@@
$x_{l,s}=-\frac{1}{2}\left(p_{l}+c^{t}_{l}+\alpha c^{o}_{l}q_{s}\frac{|S|}{\sum{q_{s}}}\right)-\frac{1}{2d}(\sum_{i\in{l^{+},l^{-}}}\mu_{i,s}a_{i,l}+\sum_{i\in{l^{+},l^{-}}}\alpha c_{i}q_{s}\frac{|S|}{\sum{q_{s}}}+\beta_{l,s})$ (8)

### v2

@@ 3 Proposed mobility hub platform design methodology
We first present the multimodal network structure and service assumptions. We then present the model formulations in detail. The model formulation for MH platform design takes the general stochastic MaaS assignment game from (Liu et al., 2025) and re-formulates it within an equivalent PURC framework with explicit MH decision variables. The goal is to provide a scalable framework that jointly decides the service decisions from the operator side and path choices from the customer side, while the platform sets service subsidies.
3.1 Network structure and model assumptions
@@
The general notation for MaaS platform design used for the MH platform design follows a multicommodity network design (MCND) problem. The multimodal network $G$, denoted by $(N,A)$, consists of all nodes $N$ and links $A$, serving all origin-destination (OD) and segment groups $S$. Origins $O$ and destinations $D$ are aggregated into centroids $\{O\cup D\}\subseteq N_{S}\subset N$. The services include a single fixed route transit (FT) operator acting as the platform for designing the MHs, and a set of MOD services that can choose to participate as feeders for these MHs through the platform. The model can be trivially modified to include fixed route buses run by other operators as feeders as well, but for simplicity in notation we assume this MH system deals only with a single fixed route platform/operator and one or more MOD feeder operators. The FT operator $f$ belongs to a singleton set $F$ and provides service on its own subnetwork $G_{f}$, which consists of node set $N_{f}\subseteq N_{F}\subset N$ and link set $A_{f}\subseteq A_{F}\subset A$. Similarly, each MOD fleet $m$ belongs to the MOD operator set $M$, with each operating a subnetwork $G_{m}$ consisting of node set $N_{m}\subseteq N_{M}\subset N$ and service link set $A_{m}\subseteq A_{M}\subset A$. A MOD service link represents a pickup at origin and drop-off at a MH, where the link attributes reflect average performance of the fleet serving this request under macroscopic modeling assumptions. Congestion effects, such as waiting for service, are captured by a node capacity at the origin. These ”in-platform” links are store-and-forward links with service queues. Each MOD operator $m$ determines the capacity for nodes $i\in N_{m}$, which are modeled as capacitated virtual links $A_{i}^{\sim}$ to access the MOD subnetwork. The capacities are used to capture queue delays that arise from congestion while waiting for service. By representing node capacities as virtual links, we maintain a link-based network structure, ensuring consistency when estimating coefficients as per Fosgerau et al. (2022).
@@
A dummy subnetwork is used to represent travel outside of the MH platform’s multimodal services (e.g., private car, telecommuting, or an alternative platform), i.e. ”out-of-platform” links. For all traveler groups $S$, a set of platform access and egress links $A_{S}^{+}$ and $A_{S}^{-}$ are used to connect respective origins and destinations with the MH based platform. Fig. 1 illustrates the multimodal network structure for one OD pair. Travelers choose their preferred paths encompassing combinations of services across the network for each OD pair segment $s\in S$, which may use the platform to access the hub and taken the fixed route transit to the destination (and vice versa), use their own modes to get to the transit station to go to the destination, or use their own mode entirely to get to the destination.
@@
In the case of a mobility hub-oriented system, we consider a FT operator also serving as a platform regulator, working with MOD operators to provide services at designated MHs through a subset of nodes $H\subset N$. These operators may be micromobility providers, ridehail, or microtransit services. The hubs in this setting represent cyberphysical gateways in which participating MOD operators that pick up or drop off passengers purchasing trip bundles within these geofenced hubs could be subsidized. We introduce MOD feeder services centered at each hub $h\in H$, with direct access to surrounding nodes $i\in N_{M}$ represented by links $A_{i,h}^{+}$. The platform operator regulates total inbound flow to ensure it does not exceed hub capacity. For consistency, these hub capacities are modeled as virtual links $A_{h}^{\sim}$.
For travelers using the hub centered platform services, an OD-based access fee is charged at the access link $A_{s}^{+}$ for $s\in S$. After paying the access fee, traversing links belonging to the platform does not incur additional ticket charges as long as the traveler does not exit the platform-based service during the trip.
Figure 1: Network illustration for one OD pair
@@
As shown in Fig. 1, we separate subnetworks to be ”in-platform” and ”out-of-platform”. We focus on the OD-based pricing at the access link for travelers and service strategy decisions of operators in-platform, including operator dedicated subsidies and service capacities. Since out-of-platform includes all other options out there (including driving, walking, telecommuting, or using MOD services via pay-as-you-go option), they are assumed to be uncapacitated and we do not explicitly model competition between the designed platform and other external modes. Instead, we focus on the cooperative game between operators in-platform.
@@
The in-platform services can be provided by multiple operators. Travelers can choose their preferred service links after entering the in-platform subnetwork by paying the access price, which is a decision variable in the model. While inside the subnetwork, only non-fee related costs are induced. For participating operators, their operating costs are covered by the subsidy provided by the platform. Other decision variables include the service capacities at either the MH or the MOD service nodes within the in-platform subnetworks. Table 1 provides the full list of notations. In addition, we list the model assumptions as follows. To avoid confusion, we define travel cost as the non-monetary penalties associated with links such as travel time. Access price is the monetary values required for entering the in-platform subnetwork. Both terms are converted to utilities.
@@
$A_{0}$ Set of transfer links between centroids and service nodes, and between layers
@@
$A_{H}$ Set of added links connecting hubs and service nodes
$A_{h}^{\sim}$ Set of virtual hub links $h\in H$ with capacity
@@
$A_{i}^{\sim}$ Set of virtual service node links $i\in N_{M}$ with capacity
$A_{i,h}^{+}$ Set of direct links from MOD node $i\in N_{M}$ to hub node $h$
$A_{S}^{+}$ Set of access links entering MH based in-platform subnetwork
@@
$z_{l}$ Capacity of access links towards MOD nodes $l\in A_{i}^{\sim},i\in N_{M}$
@@
$a_{i,l}$ Node-link incidence: 1-1 if node $i$ is origin of link $l$, +1+1 if destination, 0 otherwise
$\hat{p}_{l}$ Cap of access price for link $l\in A_{S}^{+}$
@@
$\alpha_{1}$ Weight of traveler utility
$\alpha_{2}$ Weight of operator utility
@@
$x_{l,s}$ Normalized flow on link $l\in A$ for OD pair $s\in S$
$v_{l}$ Proportion of maximum capacity opened for access link $l\in A_{0i}$
$p_{l,s}$ Service price on platform access link $l\in A_{s}^{+}$ for each OD pair $s\in S$
$r_{l}$ Operator subsidy provided by platform based on link $l\in A_{i}^{\sim}$
$b_{l}$ Capacity of hub virtual link $l\in A_{h}^{\sim}$
@@
    Travel cost of the in-platform links, FT service links, and out-of-platform links are fixed.
@@
    Transfer links between centroids and FT nodes have fixed travel cost including wait and walk time, which represent the overall cost of accessing FT services.
@@
    MOD access links are store-and-forward links with service queues, meaning that the access flow remains uncongested until capacity is reached. Once capacity is exceeded, excess flows are diverted to other links with available capacity. The queue delay is quantified by the capacity Lagrange multiplier.
@@
    All FT service links and links in the out-of-service subnetwork are uncapacitated. Congestion from background traffic are directly incorporated into the travel costs. If an out-of-service link reflects multiple modes, its travel cost reflects the expected minimized disutility option of those modes. Alternatively, each parallel mode may be modeled separately.
@@
    Operating costs of MOD services consist of service link costs and access capacity allocation costs.
  • 10.
    Operator costs for service links, MOD nodes, and mobility hubs scale linearly with demand. We acknowledge that nonlinear functions, such as the Cobb–Douglas form (Zhang et al., 2020), are often used to capture operation cost in mobility services. However, given that this model addresses macroscopic system planning under steady-state conditions, the linear assumption provides a necessary balance between model accuracy and computational efficiency.
3.2 Lower-level traveler-operator coalition choice assignment model using PURC based structure
The MH platform design problem is adapted from the stochastic MaaS assignment game model in Liu et al. (2025). The assignment game is decomposed into a bilevel problem where the upper level involves platform decisions, such as pricing and subsidies, while the lower level involves the joint decisions of the travelers and mobility operators. In the adaptation to the MH platform design problem, the platform makes decisions about how much price to charge travelers for each OD segment and how much to subsidize the MOD operators in the upper level, while the lower level determines the MH and MOD capacities and traveler flows.
Instead of using the path-based SUE-style formulation in Liu et al. (2025), we adapt the link-based PURC model proposed by Fosgerau et al. (2022) to capture the probabilistic matching mechanism. Eq. (1) is the link-based stochastic assignment problem $L_{1}$ based on the PURC concept for a MH platform.
$L_{1}\min_{\{x_{l,s},\,v_{l}\}}\quad$ $\Phi_{1}=\sum_{l\in A}\sum_{s\in S}\bar{q}d_{l}x_{l,s}^{2}$
$\quad+\alpha_{1}\bar{q}\sum_{s\in S}\left(\sum_{l\in A_{s}^{+}}d_{l}p_{l,s}x_{l,s}+\sum_{l\in A/A_{s}^{+}}d_{l}c_{l}x_{l,s}\right)$
$\quad+\alpha_{2}\left(\sum_{l\in A}\sum_{s\in S}d_{l}c^{o}_{l}x_{l,s}q_{s}+\sum_{l\in A_{i}^{\sim},i\in N_{M}}\left(z_{l}c_{i}v_{l}-\sum_{s}r_{l}q_{s}x_{l,s}\right)\right)$ (1a)
s.t. $\sum_{l\in A}a_{i,l}x_{l,s}=\begin{cases}-1,&\forall i=o,\\ 1,&\forall i=d,\\ 0,&\forall i\in N\setminus\{o,d\}\end{cases}\forall s=(o,d)\in S,\,\forall i\in N$ (1b)
$\sum_{s\in S}x_{l,s}q_{s}\leq z_{l}v_{l}\quad\forall l\in A_{i}^{\sim},\forall i\in N_{M}$ (1c)
$\sum_{l\in A^{+}_{i,h}}\sum_{s\in S}x_{l,s}q_{s}\leq b_{\tilde{l}}\quad\forall\tilde{l}\in A_{h}^{\sim},\forall h\in H$ (1d)
@@ 3.2 Lower-level traveler-operator coalition choice assignment model using PURC based structure
The decision variables are the percent link flow assignment $x_{l,s}$ and percent of maximum MOD service capacity $v_{l}$ assigned to the access links. The platform access price $p_{l,s}$ is the decision variable in the upper level model, which is presented in the following subsection. In $\Phi_{1}$, $p_{l,s}$ is treated as an input. Eq. (1b) ensures the link flow conservation. For each OD pair $s=(o,d)$ with $o$ being the origin and $d$ being the destination, the source node $o$ initiates all normalized flow, and the sink node $d$ terminates all normalized flow. Eq. (1c) controls the capacity and congestion effect of the platform service node. All access links flowing into the service nodes are bounded by the assigned capacity the operator determines. Eq. (1d) is dedicated to mobility hub facilities. For all service flows ending at mobility hubs, the sum of flow needs to be capped by the assigned mobility hub capacity, which is determined by the platform in the upper level model. Eqs. (1e - 1f) are the bounds for the two sets of decision variables.
The objective consists of three terms: the perturbed utility, traveler utility, and the operator utility. Introduced by Fosgerau et al. (2022), the PURC model follows the general RUM structure with link-based property to capture the dispersion of flows across a network. The flow assignment variable $x_{l,s}$ represents the probability of choosing link $l$ for travelers in OD pair $s$. The first term in Eq. (1a) depicts the dispersion effect. Without this term, the model becomes an all-or-nothing assignment problem. By adding a strictly increasing function where the first and second derivatives becomes zero at the origin, a dispersion effect is added that permits deviation in route choice. Fosgerau et al. (2022) demonstrate that both quadratic and entropic functions are effective. We select the quadratic form for this study due to its simplicity, and it maintains a high level of accuracy as shown in Fosgerau et al. (2022). The second term is the sum of traveler utility, which depicts the sum of all link utilities. We simplify the term by only involving the platform access price element and general travel cost element. It can be expanded to involve all utility related elements such as externalities, as long as they can be converted to utility values. For a comprehensive explanation of the disperse effect and the utility features that can be included, readers can refer to Fosgerau et al. (2022) for more information. The original PURC objective is constructed by combining the first and second terms. We show this in Proposition 1.
@@
The first and second terms combined in Eq. (1a) are equivalent to the PURC objective proposed by Fosgerau et al. (2022).
@@
$\min_{\{x_{l,s},\,v_{l},\}}\quad\Phi_{1}=\bar{q}\left(\sum_{l\in A}\sum_{s\in S}d_{l}x_{l,s}^{2}+\alpha_{1}\sum_{s\in S}\left(\sum_{l\in A_{s}^{+}}d_{l}p_{l,s}x_{l,s}+\sum_{l\in A/A_{s}^{+}}d_{l}c_{l}x_{l,s}\right)\right)$ (2)
The coefficient $\bar{q}$ is a constant. The quadratic term is the perturbed term $F(x)$, and the linear term is the link-additive utility term $U(x)$. Both terms are weighted by traversed link length $d_{l}$. The structure is therefore identical to the PURC objective proposed in Fosgerau et al. (2022) by multiplying with a constant $\bar{q}$. ∎
We introduce the third term to form the coalition between travelers and operators in a similar fashion to the model proposed by Liu et al. (2025). In Proposition 2, we show that $\Phi_{1}$ provides a flow assignment model with a traveler-operator coalition under the rule of PURC.
@@
We divide Eq. (1a) by the constant term $\bar{q}$ and write it as Eq. (3).
$\min_{\{x_{l,s},\,v_{l}\}}\quad\Phi^{\prime}_{1}={}$ $\sum_{l\in A}\sum_{s\in S}d_{l}x_{l,s}^{2}$ (3)
$+\alpha_{1}\sum_{s\in S}\left(\sum_{l\in A_{s}^{+}}d_{l}p_{l,s}x_{l,s}+\sum_{l\in A\setminus A_{s}^{+}}d_{l}c_{l}x_{l,s}\right)$
$+\alpha_{2}\left(\sum_{l\in A}\sum_{s\in S}d_{l}c^{o}_{l}x_{l,s}\frac{q_{s}}{\bar{q}}+\sum_{l\in A_{i}^{\sim}}\frac{z_{l}c_{i}v_{l}-\sum_{s}r_{l}q_{s}x_{l,s}}{\bar{q}}\right)$
@@
$\begin{split}\min_{\{x_{l,s},\,v_{i},\}}\quad L^{\prime}_{1}={}&\sum_{l\in A}\sum_{s\in S}d_{l}x_{l,s}^{2}\\ &+\alpha_{1}\sum_{s\in S}\left(\sum_{l\in A_{s}^{+}}d_{l}p_{l,s}x_{l,s}+\sum_{l\in A\setminus A_{s}^{+}}d_{l}c_{l}x_{l,s}\right)\\ &+\alpha_{2}\left(\sum_{l\in A}\sum_{s\in S}d_{l}c^{o}_{l}x_{l,s}\frac{q_{s}}{\bar{q}}+\sum_{l\in A_{i}^{\sim}}\frac{z_{l}c_{i}v_{l}-\sum_{s}r_{l}q_{s}x_{l,s}}{\bar{q}}\right)\\ &+\sum_{s\in S}\sum_{i\in N}\mu_{i,s}(\sum_{l}a_{i,l}x_{l,s}-f_{i,s})\\ &+\sum_{l\in A_{i}^{\sim},i\in N_{M}}\lambda_{l}(\sum_{s\in S}x_{l,s}q_{s}-z_{l}v_{l})\\ &+\sum_{\tilde{l}\in A_{h}^{\sim},h\in H}\varphi_{\tilde{l}}(\sum_{l\in A_{i,h}^{+}}\sum_{s\in S}x_{l,s}q_{s}-b_{\tilde{l}})\\ &+\sum_{l\in A}\sum_{s\in S}\beta_{l,s}(x_{l,s}-1)+\sum_{l\in A_{i}^{\sim},i\in N_{M}}\pi_{l}(v_{l}-1)\end{split}$ (4)
@@
$\begin{split}\frac{\partial L^{\prime}_{1}}{x_{l,s}}&=2d_{l}x_{l,s}+d_{l}\left(\alpha_{1}\left(p_{l}+c_{l}\right)+\alpha_{2}\frac{(c^{o}_{l}-r_{l})q_{s}}{\bar{q}}\right)\\ &+\sum_{i\in{l^{+},l^{-}}}\mu_{i,s}a_{i,l}+\lambda_{l}q_{s}+\sum_{\tilde{l}^{-}\in l^{+}}\varphi_{\tilde{l}}q_{s}+\beta_{l,s}=0\end{split}$ (5)
We slightly change the notation of the term associated with $\alpha_{1}$ in Eq. 5 for simplicity. By setting $c_{l}$ to zero when $l\in A_{s}^{+}$ and $p_{l}$ to zero when $l\in A\setminus A_{s}^{+}$, Eq. 5 is equivalent to the first order derivative of Eq 4. Because the chosen links will have strictly positive flows, the $v_{l}$ associated with those links also fulfill the optimality conditions.
$\frac{\partial L^{\prime}_{1}}{v_{l}}=\alpha_{2}\frac{z_{l}c_{i}}{\bar{q}}-\lambda_{l}z_{l}+\pi_{l}=0$ (6)
By re-writing Eq. (6), we obtain $\lambda_{l}=\frac{\alpha_{2}c_{i}}{\bar{q}}+\frac{\pi_{l}}{z_{l}}$. By plugging it in Eq. (5), we have Eq. (7).
$\frac{\partial L^{\prime}_{1}}{x_{l,s}}$ $=2d_{l}x_{l,s}+d_{l}\left(\alpha_{1}\left(p_{l}+c_{l}\right)+\alpha_{2}\frac{(c^{o}_{l}-r_{l})q_{s}}{\bar{q}}\right)$ (7a)
$+\sum_{i\in{l^{+},l^{-}}}\mu_{i,s}a_{i,l}+\sum_{i\in l^{+}}\alpha_{2}c_{i}\frac{q_{s}}{\bar{q}}+\frac{\pi_{l}q_{s}}{z_{l}}+\sum_{l^{\prime-}\in l^{+}}\varphi_{l^{\prime}}q_{s}+\beta_{l,s}=0$ (7b)
@@
$x_{l,s}=-\frac{1}{2}\left(\alpha_{1}\left(p_{l}+c_{l}\right)+\alpha_{2}\frac{(c^{o}_{l}-r_{l})q_{s}}{\bar{q}}\right)-\frac{1}{2d}(\sum_{i\in{l^{+},l^{-}}}\mu_{i,s}a_{i,l}+\sum_{i\in l^{+}}\alpha_{2}c_{i}\frac{q_{s}}{\bar{q}}+\frac{\pi_{l}q_{s}}{z_{l}}+\sum_{l^{\prime-}\in l^{+}}\varphi_{l^{\prime}}q_{s}+\beta_{l,s})$ (8)